OpenStudy (anonymous):

how do you solve 2+4x+x square=0 and getting the answers -0.59, -3.41

OpenStudy (amistre64):

combine like terms: 2 +4x +x = 0.....x square?

OpenStudy (amistre64):

x^2 + 4x + 2 = 0 -2 +- sqrt(2)

OpenStudy (anonymous):

yes x square i dont know how to write the little two on the top

OpenStudy (amistre64):

the 6 button has a caret if your shift it :) x^2

OpenStudy (anonymous):

ok.. so how do i get the answers -0.59 and -3.41 by doing this equation???

OpenStudy (anonymous):

\[x^2+4x+2=0\] is a quadratic equation ( a polynomial of second degree).. The discriminant is \[Δ=4^2-4*1*2=8\] so the roots are \[x_1=(-4+\sqrt{Δ})/(2*1)=-0.59\] and \[x_2=(-4-\sqrt{Δ})/(2*1)=-3.41\] For more information try: http://en.wikipedia.org/wiki/Quadratic_equation

OpenStudy (anonymous):

I should add that when you have a quadratic equation \[ax^2+bx+c=0\] the discriminant is \[Δ=b^2-4ac\] and the roots are \[x1=(-b+\sqrt{Δ})/2a\] and \[x2=(-b-\sqrt{Δ})/2a\]